# MolBind: Multimodal Alignment of Language, Molecules, and Proteins

### v1

@@ 2.3 Dataset Collection
In this section, we briefly describe how to construct the MolBind-M4 dataset (full details in Appendix A.2), including 325K pairs of language-graph data, 161K pairs of language-conformation data, 161K pairs of graph-conformation data, and 72K pairs of conformation-protein data.
Language-Graph Pairs: PubChem Kim et al. (2023), the most extensive source of chemical information, provided many molecules with diverse descriptions. We transformed these molecules into graphs and paired them with corresponding textual descriptions, yielding 32K language-graph pairs.
Language-Conformation Pairs: Using Molecule3D Xu et al. (2023), which contains 3.9M ground-state molecular conformations, and the GEOM dataset Axelrod et al. (2022) with 37M higher-quality conformations, we matched molecular IDs (CIDs) and InChIs in them with textual descriptions from PubChem. This process resulted in 157K language-conformation pairs.
@@
Graph-Conformation Pairs: Converting 3D molecular structures from the language-conformation dataset into 2D graphs allowed us to generate 158K graph-conformation pairs.
@@ A.2 Details of Constructing MolBind-M4
Language-Graph Pairs. We used Pug View Kim et al. (2019) to download molecule annotations. In total, we obtained 349,356 molecules with textual descriptions. Most of these annotations begin with the common name or the International Union of Pure and Applied Chemistry (IUPAC) name Kuhn et al. (2004). To encourage the text encoder to learn generalized descriptions of molecule functionality and properties, we uniformly replaced descriptions that start with a name with ’The molecule.’ We sampled high-quality validation and test sets, each consisting of 1,500 pairs with texts longer than 19 words. The remaining samples were allocated to a pre-training set, totaling 319,353 samples.
@@
Specifically, we initially acquired molecular conformations from two sources. The first dataset, Molecule3D Xu et al. (2021b), includes 39M ground-state molecular conformations computed via DFT from PubChemQC Nakata and Shimazaki (2017). These conformations were then matched with corresponding textual descriptions in language-graph pairs using CID as a unique identifier, resulting in 25, 961 conformation-language pairs. The second dataset, GEOM, contains approximately 37M conformations generated using a semi-empirical DFT algorithm. Due to the absence of CIDs, we matched conformation and language-graph pairs using the International Chemical Identifier (InChI) Heller et al. (2015) to obtain textual descriptions and identify molecules. Through this process, we obtained a total of 158, 327 high-quality conformations. For data partitioning, we sampled 1,500 samples for both the training and testing sets from the most accurate ground-state conformations derived from Molecule3D. During the sampling process, we ensured that the text did not appear in the training set of language-graph pairs to mitigate potential data leakage.

### v2

@@ 2.3 Dataset Collection
In this section, we briefly describe how to construct the MolBind-M4 dataset (full details in Appendix A.2), including 322K pairs of language-graph data, 161K pairs of language-conformation data, 161K pairs of graph-conformation data, and 72K pairs of conformation-protein data.
Language-Graph Pairs: PubChem Kim et al. (2023), the most extensive source of chemical information, provided many molecules with diverse descriptions. We transformed these molecules into graphs and paired them with corresponding textual descriptions, yielding 322K language-graph pairs.
Language-Conformation Pairs: Using Molecule3D Xu et al. (2023), which contains 3.9M ground-state molecular conformations, and the GEOM dataset Axelrod et al. (2022) with 37M higher-quality conformations, we matched molecular IDs (CIDs) and InChIs in them with textual descriptions from PubChem. This process resulted in 161K language-conformation pairs.
@@
Graph-Conformation Pairs: Converting 3D molecular structures from the language-conformation dataset into 2D graphs allowed us to generate 161K graph-conformation pairs.
@@ A.2 Details of Constructing MolBind-M4
Language-Graph Pairs. We used Pug View Kim et al. (2019) to download molecule annotations. In total, we obtained 322,353 molecules with textual descriptions. Most of these annotations begin with the common name or the International Union of Pure and Applied Chemistry (IUPAC) name Kuhn et al. (2004). To encourage the text encoder to learn generalized descriptions of molecule functionality and properties, we uniformly replaced descriptions that start with a name with ’The molecule.’ We sampled high-quality validation and test sets, each consisting of 1,500 pairs with texts longer than 19 words. The remaining samples were allocated to a pre-training set, totaling 319,353 samples.
@@
Specifically, we initially acquired molecular conformations from two sources. The first dataset, Molecule3D Xu et al. (2021b), includes 39M ground-state molecular conformations computed via DFT from PubChemQC Nakata and Shimazaki (2017). These conformations were then matched with corresponding textual descriptions in language-graph pairs using CID as a unique identifier, resulting in 25, 961 conformation-language pairs. The second dataset, GEOM, contains approximately 37M conformations generated using a semi-empirical DFT algorithm. Due to the absence of CIDs, we matched conformation and language-graph pairs using the International Chemical Identifier (InChI) Heller et al. (2015) to obtain textual descriptions and identify molecules. Through this process, we obtained a total of 158, 237 high-quality conformations. For data partitioning, we sampled 1,500 samples for both the training and testing sets from the most accurate ground-state conformations derived from Molecule3D. During the sampling process, we ensured that the text did not appear in the training set of language-graph pairs to mitigate potential data leakage.